Jeff is saving for his retirement 21 years from now by setting up a savings plan. He has set up a savings plan wherein he will deposit $ 149.00 at the end of each month for the next 14 years. Interest is 5 % compounded monthly.

​(a) How much money will be in his account on the date of his​ retirement?

​(b) How much will Jeff ​contribute? ​

(c) How much will be​ interest?
Answer #1

a)
Nper = 21 * 12 = 252
PMT = 149
Rate = 5% / 12
PV = 0

Future value can be calculated by using the following excel formula:
=FV(rate,nper,pmt,pv)
=FV(5%/12,252,-149,0)
= $66,206.93

Amount of money will be in his account on the date of his​ retirement = $66,206.93

b)
Amount of money jeff contribute = Number of period * Monthly deposit
= 252 * $149
= $37,548

c)

Interest = $66,206.93 - $37,548 = $28,658.93